OpenStudy (rational):

how to show \(n^3-4\) is not divisible by \(13\) for all \(n \in \mathbb{Z}\)

OpenStudy (rational):

heyy

Nnesha (nnesha):

who??? me ??

OpenStudy (rational):

green apple

Nnesha (nnesha):

yes sSIR ??

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

*

Nnesha (nnesha):

^^^ what does that mean ??

OpenStudy (rational):

few observations : 1) \(n^{12}-4\) is not divisible by \(13\) because \(n^{12}\) leaves a remainder \(1\) or \(0\) by fermat's little theorem 2) first observation implies \(n^{6}\) has to leave a remainder \(-1\) or \(0\)

Nnesha (nnesha):

is it \[n^3 \] or \[n^{12}\] ??

OpenStudy (rational):

Ahh our goal is to show n^3 - 4 is not divisible by 13, was just messing with higher powers because they are easy

Nnesha (nnesha):

hmmmm nvm ^_^ srry

OpenStudy (rational):

One straight forward way is like this : Every integer can be expressed in one of the forms {13k, 13k+1, 13k+2, ... , 13k+12} It is sufficient to show that the given statement is true for these 13 cases

OpenStudy (rational):

case1 : n = 13k \(n^3 - 4 = (13k)^3 - 4 \equiv -4 \pmod{13}\) the remainder is -4, so the statement is true for numbers of form 13k case2 : n = 13k+1 \(n^3 - 4 = (13k+1)^3 - 4 \equiv 1-4\equiv -3 \pmod{13}\) the remainder is -3, so the statement is true for numbers of form 13k+1 case3 : n = 13k+2 \(n^3 - 4 = (13k+2)^3 - 4 \equiv 8-4\equiv 4 \pmod{13}\) the remainder is 4, so the statement is true for numbers of form 13k+2 etc... this works, but im looking for a more neat/compact method

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (anonymous):

still need help , brb

OpenStudy (rational):

im looking for an alternative method to prove this w/o using division algorithm

OpenStudy (anonymous):

ok using this theorem

OpenStudy (anonymous):

\(n^3\equiv 4 \mod 13 \text{ has a solution }\iff 4^{\phi(13)/d} \equiv 1 \mod 13 \\ d=gcd(\phi(13),3)=3\\ \begin{align*} 4^{\phi(13)/3} &=4^{12/3} \\ &= 4^{4}\\ &\equiv 9 \mod 13\not\equiv 1 \mod 13 \end{align*} \\ \text{thus } n^3\equiv 4\mod 13 \text{ has no solution} \\ \rightarrow n^3-4\equiv 0\mod 13 \text{ has no solution}\rightarrow 13 \not| n^3-4 \)

OpenStudy (anonymous):

\(\text{ some other trick with D.A}\\ n\equiv 0,1,2,3,4,5,6,7,8,9,10,11,12 ~\mod 13\\ n^3\equiv -1,0,1,5,8 ~\mod 13\\ \text{ we need the residues that satisfied }\\ n^3-4 \equiv 0 \mod 13\\ \downarrow \\ \begin{align*} -1-4 \mod 13 \equiv 8 &\not\equiv 0 \mod 13 \\ 0-4 \mod 13 \equiv 9 &\not\equiv 0 \mod 13 \\ 1-4 \mod 13 \equiv 10 &\not\equiv 0 \mod 13 \\ 5-4 \mod 13 \equiv 1 &\not\equiv 0 \mod 13 \\ 8-4 \mod 13 \equiv 4 &\not\equiv 0 \mod 13 \end{align*} \)

OpenStudy (anonymous):

can also simply say there is no ind 4 mod 13 integer solution ?

OpenStudy (kainui):

\[\large n^3 -4 \ne 0 \mod 13\] We need to prove that this equation is never satisfied, if n is odd, then it won't be satisfied, so we just need to check even values of n. We can represent all numbers as \[\large 13k+m\] with k as any integer and we allow m to vary between 0 and 12 to cover all possible numbers. \[\large 0 \le m \le 12\] But like we said, we don't care about odd numbers, only even numbers so let's multiply 13k+m by 2 and let that be our n: \[\large n = 2(13k+m)=26k+2m\] Now let's just plug it in: \[\large (26k+2m)^3 -4 \equiv 8m^3-4 \mod 13 \\ \large \equiv 4(2m^3-1) \mod 13\] I guess now I got stuck, I don't really know enough about modular arithmetic to do more with this, I was just thinking that in order for this to work the (2m^3-1) part has to be 13 which means m^3=7 in mod 13 arithmetic, but since they're both coprime then there is no solution, which is what we want. But I'm not really sure if that reasoning works here or not, maybe someone can help.

OpenStudy (rational):

that looks neat and close to division algorithm argument ! xD We need to prove that this equation is never satisfied, `if n is odd, then it won't be satisfied, so we just need to check even values of n`. We can represent all numbers as how do we know odd n's wont work ?

OpenStudy (kainui):

well since odd*odd = odd we know n^3 is odd if n is odd. We also know odd + even = odd so since n^3 is odd, and -4 is even we still have an odd... But now I see a problem... 0 = 13 mod 13 so this argument completely is broken! haha oh well.

OpenStudy (rational):

but still it proves that there are no solutions in even integers, that part is flawless :)

OpenStudy (anonymous):

we are roaming around D.A here as will

OpenStudy (kainui):

Ok here's a new argument, just show that \[\large n^3-4 \ne 0 \mod 13\] and let \[\large n=13k+a \\ \large k \in \mathbb{Z} \\ \large 0 \le a \le 12\] When we plug it in, we have \[\large (13k+a)^3 -4 \ne 0 \\ \large a^3-4 \ne 0 \\ \large a^3 \ne 4\] then we have to somehow show that no cube of any number from 0 to 12 is 4 mod 13. So one way is to just compare the 12 perfect cubes and make sure none of them are 4 larger than a multiple of 13, but that's basically what was said already which is kind of lame.

OpenStudy (rational):

thats division algorithm and yes it works by exhausting all the 12 cases

OpenStudy (anonymous):

its what i did here xD \(\uparrow \)but worked more less

OpenStudy (kainui):

Hahaha sorry I don't know what division algorithm means exactly but I think I kinda get the idea anyways

OpenStudy (anonymous):

ok here is an idea you know according to division algorithm we can all numbers like this :- \(13k\\ 13k+1\\ 13k+2\\ 13k+3\\ 13k+4\\ 13k+5\\ 13k+6\\ 13k+7\\ 13k+8\\ 13k+9\\ 13k+10\\ 13k+11\\ 13k+12\)

OpenStudy (anonymous):

which is the same as \(13k \equiv 0 \mod 13\\ 13k+1\equiv 1 \mod 13\\ 13k+2\equiv 2 \mod 13\\ 13k+3\equiv 3 \mod 13\\ 13k+4\equiv 4 \mod 13\\ 13k+5\equiv 5 \mod 13\\ 13k+6\equiv 6 \mod 13\\ 13k+7\equiv 7 \mod 13\\ 13k+9\equiv 9 \mod 13\\ 13k+10\equiv 10 \mod 13\\ 13k+11\equiv 11 \mod 13\\ 13k+12\equiv 12 \mod 13\) so make ur self free to take mod to power 3,7 or what ever u wont to see at which modulo of 13 it follow